• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mindy

    Log entry time 00:03:03 on June 7,2010

    Entry number 325005

    keyword=06/06/10 Swing Shift Summary

    started from production but MCC called to request for an access to fix the IOC/BPM problem.

    during the downtime, Rupesh, Sasha, Romain and Jixie were doing the large charge asymmetry measurement using the parity DAQ as well as the Moller DAQ, respctively. Megan also made an halog entry to yield us a charge asymmetry from Compton DAQ.

    Since MCC took a while to fix BPM problem, we didn't have the beam available until 9:30 PM. Then, Rupesh started his pedestal test and the large charge asymmetry measurement again!

    After Rupesh's test, we just started the long beam recovery process.

    Now, the beam is back, and we are now going to the production mode again!